Using a rope, a person pulls a sled 47 m across level ground through the snow with a force of 79 N directed at an angle of 22.3o above the horizontal. How much work was done on the toboggan by the tension? (Friction is negligible.)
1
Expert's answer
2020-02-05T12:28:53-0500
"W=Fs\cos{22.3\degree}"

"W=(79)(47)\cos{22.3\degree}=3435\ J"